OpenStudy (anonymous):

If 9(x-9) = -11 then x = ? I thought I could just distribute it but the answers are in fraction form. I'm not sure how to find the fraction unless I just substitute the answers in a calculator. It's not a homework question, it's from an ACT prep book.

OpenStudy (shamil98):

Distribute the 9 to what is inside the parantheses. I'll get you started. \[\large 9(x-9) = -11\] \[\large 9x - 81 = -11\]

OpenStudy (shamil98):

Can you solve for x from here?

OpenStudy (anonymous):

All I have figured out is that x rounds up to about 7.7 in decimal form

OpenStudy (shamil98):

9x - 81 = -11 9x = 70 x = 70/9 x = 7.7 correct